OpenStudy (anonymous):

Medal Write as a fraction 66 2/3%

jagr2713 (jagr2713):

dYou do it the way you change any percentage to a fraction. Divide by 100% and simplify. 66 2/3% = (66 2/3%)/(100%) = (200/3)/100 = 200/300 = (2*100)/(3*100) = (2/3)*(100/100) what do we get?

jagr2713 (jagr2713):

@harz360

OpenStudy (anonymous):

yes i understand

OpenStudy (anonymous):

so what is the answer ?

jagr2713 (jagr2713):

I am asking you :DD

OpenStudy (anonymous):

2/3

jagr2713 (jagr2713):

Correct :D

OpenStudy (anonymous):

thank you

jagr2713 (jagr2713):

Anytime :D